Nevey:

What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 8 2x + 3y = 2 4 −2 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system

dude:

Let's solve for this algebraically, choose either the first or the second equation

Nevey:

1st

dude:

Wait not theres an easier way to do this lol

dude:

Which variable do you want to get rid of, x or y?

Nevey:

y

dude:

Right, so what are the common multiples for 2 and 3?

Nevey:

x and y?

dude:

Do you know what multiples are?

Nevey:

kinda sorta but not really

dude:

Numbers that can be divided by the number So like 10 is a multiple of 2 because \(\frac {10}{5}\) is 2 Or any whole number multiplied by 2 ¯\_(ツ)_/¯ 2,4,6,8,10,14 and so on

Nevey:

oh ok i get it

Nevey:

so 6 right?

dude:

I have you the ones for 2, can you give me for 3?

dude:

Yep

Nevey:

yay :)

dude:

Lets multiply the top equation by 3 and the bottom equation by -2 (To get them to cancel out) \(~~~~3(3x + 2y = 8)=?\\ ~\\ -2(2x + 3y = 2)=?\) Can you solve both of these? (Distribute)

Nevey:

(9x + 6y = 8) = ? (-4x + -6y = 2) = ?

dude:

Multiply the 3 and 8, the whole equation is being multiplied

Nevey:

oh ok yea

dude:

Same with the second equation

Nevey:

(9x + 6y = 24) = ? (-4x + -6y = -4) = ?

Nevey:

right?

dude:

Right, now add both equations \( ~~~~~~~~~~~~9x + 6y = 24 \\ ~~+~-4x + -6y = -4 \\ \_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\)

Nevey:

btw, i marked -2 before i asked the Q, meaning thats what i think it is. is that right?

dude:

Yeah its right

Nevey:

ok so all this wasnt even needed XD

Nevey:

ty anyway :)

dude:

Well this is the whole process if you want to know it ¯\_(ツ)_/¯

dude:

No problem

Nevey:

the next one is that last one and i didnt answer it meaning i really need help with this one..